OpenStudy (anonymous):

Find a cubic function with the given zeros. (x-sqrt2)(x+sqrt2)(x+2)

OpenStudy (tyteen4a03):

In layman's terms, expand this expression.

OpenStudy (anonymous):

\[(x^2+x \sqrt{2}-x \sqrt{2}-(\sqrt{2}*\sqrt{2}))(x+2)\]

OpenStudy (anonymous):

how do you multiply the sqrts

OpenStudy (tyteen4a03):

Remember that \((x - y)(x + y) = x^2 - y^2\).

OpenStudy (anonymous):

\[(x^2+x \sqrt{2}-x \sqrt{2}-2)(x+2)\]

OpenStudy (tyteen4a03):

@torobi Look at my formula and apply it to the first two brackets. Also, \(\sqrt{y} * \sqrt{y} = y.\)

OpenStudy (anonymous):

what if \[-\sqrt{y} * \sqrt{y}\]

OpenStudy (tyteen4a03):

in that case, it will be -y. Think \(-(\sqrt{y} * \sqrt{y})\)

OpenStudy (anonymous):

ok so my answer was is that correct